OpenStudy (anonymous):

For the simple harmonic motion described by the trig functioon, find (a) the maximum displacement, (b) the frequency, (c) the value of d when t = 5, and (d)) the least positive value of t for which d = 0. d= 1/64 sin792 π t

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

The maximum displacement is simply the amplitude of d. The frequency can be found from the relation \(\omega=2\pi f\implies f=\cfrac{\omega}{2\pi}\). From d, \(\omega=792\pi\). Now just solve for \(f\), in hertz. To find the value of d at \(t=5\), just plug into d. When d=0, the sine value is zero. For what value of t is sine=0. If zero is excluded, the smallest positive value is when \(792\pi t=2\pi\). Now just solve for t.

OpenStudy (anonymous):

so for (a) the answer is 1/64?

OpenStudy (anonymous):

(b) is 1/64?

OpenStudy (ybarrap):

a is right b is wrong, should be 792/2

OpenStudy (anonymous):

alright, so for b it would be 792/2= 396?

OpenStudy (ybarrap):

make sense?

OpenStudy (anonymous):

its 396 for (b), right?

OpenStudy (ybarrap):

792/2, whatever that is

OpenStudy (anonymous):

alrighty.

OpenStudy (anonymous):

For c what do u mean by pluggin it in.

OpenStudy (ybarrap):

$$ d= \left (1/64 \right )\sin(792 π (5)) $$

OpenStudy (anonymous):

I get 6.125^-12 on calculator...

OpenStudy (ybarrap):

That's a pretty small number, might as well be zero: \(792\times 5\) is a multiple of \(2\) and \(\sin(2n\pi)=0\) for \(n\in\mathbb N\)

OpenStudy (anonymous):

so for (c) the answer is 0?